Your company is starting a new product line that will require buying a new piece of manufacturing equipment in Year 0 for $25Please show work
Answer #1

Initial Cost = 250,000

Profit in the 1st year = 30,000, and to grow by 4% each year after that.

a. What is the simple payback period?

Cash Flow

Net Cash Flow

Year 0

$-250,000.00

$-250,000.00

Year 1

$30,000.00

$-220,000.00

Year 2

$31,200.00

$-188,800.00

Year 3

$32,448.00

$-156,352.00

Year 4

$33,745.92

$-122,606.08

Year 5

$35,095.76

$-87,510.32

Year 6

$36,499.59

$-51,010.74

Year 7

$37,959.57

$-13,051.17

Year 8

$39,477.95

$26,426.79

Simple pay-back period = 7+(13,051 ÷ 39,478) = 7.3 years

Rounding off, Pay back period = 7 years (taking integers)

b. What is discounted payback period at an interest rate of 6%.

Cash Flow

Discounted Cash Flow

Net Discounted Cash Flow

Year 0

$-250,000.00

$-250,000.00

$-250,000.00

Year 1

$30,000.00

$28,301.89

$-221,698.11

Year 2

$31,200.00

$27,767.89

$-193,930.22

Year 3

$32,448.00

$27,243.97

$-166,686.26

Year 4

$33,745.92

$26,729.93

$-139,956.33

Year 5

$35,095.76

$26,225.59

$-113,730.74

Year 6

$36,499.59

$25,730.77

$-87,999.97

Year 7

$37,959.57

$25,245.28

$-62,754.69

Year 8

$39,477.95

$24,768.96

$-37,985.73

Year 9

$41,057.07

$24,301.62

$-13,684.11

Year 10

$42,699.35

$23,843.10

$10,158.98

Discounted payback period = 9+[-13684 – 0 ÷ -13684 – (10159)]*1% = 9.5 years

Rounding off, payback period = 10 years (taking integers)

c) If the company looks for a discounted payback period of 6 years, is it a good investment?

B. No. At a discounted payback period of 6 years, the investment is not good. The investment should be rejected as both the simple payback period and discounted payback period is more than the standard given.